# Quasiperiodicity protects quantized transport in disordered systems without gaps

### v1

@@ Abstract
The robustness of topological properties, such as adiabatic quantized currents, generally depends on the existence of gaps surrounding the relevant energy levels or on symmetry-forbidden transitions. We uncover a new mechanism for topological protection via the observation of quantized currents that survive the addition of bounded local disorder beyond the closing of the relevant instantaneous energy gaps in a driven Aubry-André-Harper chain, a prototypical model of quasiperiodic systems. Using a local picture in configuration-space based on Landau-Zener transitions, we show that in this case the topological protection does not depend on the spectral gaps, but rather on the number of states in an occupied band and the position of the band in configuration space. Moreover, we propose a protocol, directly realizable in for instance cold atoms or photonic experiments, which leverages this stability to prepare topological many-body states with high Chern numbers and opens new experimental avenues for the study of both the integer and fractional quantum Hall effects.
@@
A key feature of topological phases is the robustness to local perturbations that leads to quantized physical observables characterized by a bulk topological invariant, a famous example of which is found in the integer quantum Hall effect [1, 2]. Topological phases can also be realised in dynamical systems, where time is interpreted as an additional virtual dimension. The canonical example is Thouless pumping [3, 4], where slow time-periodic modulation of a typically 1D Hamiltonian generates quantized charge transport indexed by virtual 2D Chern numbers [5, 6]. Experimentally, Thouless pumps have been realized within various platforms, e.g. photonic waveguides [7] or ultracold atoms [8, 9, 10, 11]. They are typically considered in the context of periodic systems, where the pumped charge is only quantised when the pumping time is an exact multiple of the pump period. In contrast, Thouless pumping in quasiperiodic systems can exhibit constant quantized pumped charge and currents at all times during the pump cycle [12].
@@
The stability of quantized Thouless pumping to disorder remains a topic of active research [4, 13, 14]. More generally, it has long been shown mathematically that the adiabatic theorem does not require a spectral gap [15], even in open quantum systems [16]. For experimentally accessible observables such as the pumped charge, on the other hand, topological quantisation typically remains stable to generic perturbations only up to the closing of spectral gaps or mobility gaps – gaps between delocalized eigenstates [17] – and/or the breaking of explicit symmetry protections [18, 19, 20, 21]. In particular, it was shown that Thouless pumps in periodic systems break down under closing of the energy gaps by disorder [22, 23, 24]. In this letter, we report on an experimentally accessible Thouless pump for quasiperiodic systems, characterized by persistent quantized pumping in the presence of large bounded local disorder, surviving the disorder-induced closure of spectral gaps. The topological protection does not rest on any spectral condition or symmetry protection and instead depends on the number of states in the pumped band, in stark contrast with previously known examples.
@@
To explain this counter-intuitive robustness, we develop a perturbative argument based on local Landau-Zener (avoided-crossing) transitions, and derive analytical expressions for the critical disorder strengths. The enhanced stability results from two special properties of 1D quasiperiodic Hamiltonians: first, Anderson localization [25] of the eigenstates even in the absence of additional disorder, and second, the bounded Cantor set spectrum of quasiperiodic operators – a closed no-where dense set for which any element itself is a limit point with a fractal hierarchy of gaps [26] – whose integrated density of states (IDoS) is quantized by the gap labeling theorem [27, 28, 29]. Further, we discuss how the robustness can be readily probed within existing cold atoms experiments, and leveraged to prepare many-body states with arbitrary Chern number.
@@
While our results generalize to other short-range, bounded, analytic quasiperiodic Hamiltonians, [^1] Our results rest on the gap labeling theorem and Anderson localization which hold for this broad class of quasiperiodic models [27, 25]. we focus on the prototypical Aubry-Andre-Harper (AAH) model as realized in e.g. optical lattices [30, 31], photonic systems [5], or with cavity polaritons [32] and known as the almost Mathieu operator in the mathematical physics literature
$\hat{H}_{AAH}^{J,V}(\varphi)=\hat{V}(\varphi)-\hat{J},$ (1)
consisting only of a nearest-neighbor hopping term
$\hat{J}=\sum_{x}J(\hat{a}^{\dagger}_{x+1}\hat{a}_{x}+\hat{a}^{\dagger}_{x}\hat{a}_{x+1})$ (2)
and an on-site cosine potential
$\hat{V}(\varphi)=\sum_{x}2V\cos(2\pi\beta x+\varphi)\hat{a}^{\dagger}_{x}\hat{a}_{x},$ (3)
@@
Figure 1: Perturbative derivation of the AAH energy spectrum: (a) AAH chain where grey line denotes the quasiperiodic modulation of onsite energies $\epsilon_{i}$, and circled dots are lattice sites. Solid lines indicate sites connected by the tunneling $\hat{J}$. Resonances between $n$-th nearest neighbors (blue ellipses show the $n=1$ resonances) are mediated by an exponentially decreasing effective tunneling rate $J_{eff}^{(n)}\propto J^{n}/V^{n-1}$. (b) We map the AAH chain onto a uniformly dense unit circle – referred to as configuration space. In this representation, nearest-neighbors are separated by an angle $\theta_{i+1}-\theta_{i}=2\pi\beta$, and the resonances occur at fixed angles $\theta^{(n)}$ (blue lines show $n=1$ resonances). (c) The resonances result in a dense set of infinitely many pairs of energy gaps labelled by the integer $n$, which for $V\gg J$ lie at energies $E^{(\pm n)}$ (we show gaps up to $n=3$, but all gaps are open for all values of $V/J$). Sets of states surrounded by two selected gaps acquire a Chern number fixed by the labels of the surrounding gaps (we show the Chern numbers of bands up to $n=3$). Parameters: $\beta=\sqrt{2}/2$.
@@ I AAH: Energy gaps and configuration space
In periodic systems, the Brillouin zone offers a suitable base manifold on which to calculate topological invariants and band structures. To circumvent the lack of a Brillouin zone for quasiperiodic lattices, we construct a configuration space representation of the clean AAH chain, see Fig. 1 b. This representation offers a powerful language to discuss such quasiperiodic systems and enables a simple perturbative argument to derive the positions of the energy gaps, their labels, and exact expressions for their respective integrated densities of state. It is based on the observation that the only degree of freedom that differentiates one lattice site from another is the local value of the phase of the quasiperiodic modulation in Eq. 3. Therefore, labeling the sites with the local phase of the modulation offers a convenient description of the problem. In this configuration space representation, the position of site $i$ is mapped to the angle $\theta_{i}$ on the unit circle according to
@@
see Fig. 1b. Since $\beta$ is irrational, this map is ergodic and the set of all possible $\theta_{i}$, $\{\theta_{i}\}_{i\in\mathbb{Z}}$, densely and uniformly populates the unit circle, implying that every site is identified by a unique value of $\theta$.
@@
is dense on the interval $\left[-2V,2V\right]$ (see right-hand side of Fig. 1 c), and every eigenstate is localized on a single lattice site. Turning on a small tunneling amplitude ($J/V\ll 1$), nearest neighbors whose onsite energy difference is small $\delta E=\epsilon_{i+1}-\epsilon_{i}\ll J$ (referred to as resonant neighbors, marked in blue in Fig. 1 a) hybridize and form (anti-) symmetric dimers, separated by the energy $\Delta E^{(1)}=2J$. In addition, nearly resonant neighbors with similar onsite energies will also be affected by this hybridisation-induced level repulsion.
@@
which causes the opening of a pair of gaps of size $\Delta E^{(n)}\propto 2J^{n}/V^{n-1}$, at energies
@@
Here, $[n\beta]$ refers to the integer part of $n\beta$. Eq. 9 shows that the AAH spectrum contains an infinite hierarchy of pairs of gaps – forming a dense set – labelled by the integer Chern numbers ±$\pm n$, where the two branches of the equation ensure consistency with the direction of quantized currents (see Appendix A for details). By the gap labelling theorem, every band, i.e., the set containing all states between two selected energy gaps, is indexed by a virtual Chern number [6] fixed by the difference between the two surrounding gap labels (Fig. 1 c). This Chern number is referred to as “virtual” in the sense that it is defined by treating $\varphi$ as a virtual dimension of the system. It arises from the non-commutativity of $x,\partial_{\varphi}$-translations in configuration space analogous to the non-commutativity of $x,y$-translations responsible for the integer quantum hall effect [6, 35, 36].
While the above gap construction relies on the hierarchy of effective tunnelings $J_{\textit{eff}}^{(n+1)}\ll J_{\textit{eff}}^{(n)}$ to generate a hierarchy of spectral gaps, as mentioned above, the spectrum indeed forms a no-where dense Cantor set and the IDoS is constant for all $V/J\neq 1$ [26, 28], extending the results of the above analysis to any value of $V/J>1$ and by self-duality to $V/J<1$.
@@ II Thouless Pumping in the AAH model
The Hamiltonian $\hat{H}^{J,V}_{AAH}(\varphi)$ is smooth and periodic in $\varphi$ and, as discussed before, its spectrum forms a Cantor set independent of $\varphi$ whose gaps are labeled by Chern numbers [26]. Since a Cantor set is nowhere dense, the spectrum is separated into arbitrarily many “bands”
@@
isolated by spectral gaps $\Delta E^{(m)},\Delta E^{(n)}$ at energies $E^{(m)},E^{(n)}$ independent of $\varphi\in[0,2\pi)$. This guarantees the existence of smooth contours $\gamma_{m,n}(\varphi)$ in the complex energy plane containing all of $B^{(m,n)}$ for all $\varphi$, which in turn guarantees adiabatic transport in the $\dot{\varphi}\rightarrow 0$ limit [37]. Thus, the 1D system evolving under $\hat{H}^{J,V}_{AAH}(\varphi(t))$ with $\varphi(t)=2\pi\dot{\varphi}t+\varphi_{0}$ and $(\hbar 2V\dot{\varphi})^{1/2}\ll\min(\Delta E^{(m)},\Delta E^{(n)})$ constitutes a Thouless pump and results in a quantized current proportional to the band’s Chern number. The current
$\hat{I}(t)=\frac{i}{L}J\sum_{x=1}^{L}(\hat{a}^{\dagger}_{x+1}\hat{a}_{x}-\hat{a}^{\dagger}_{x}\hat{a}_{x+1})$ (11)
@@
is the spatial average of the current operator over the system (of length $L$), and is computed as the derivative of the total pumped charge $I\equiv\partial_{\varphi}Q$ averaged over the entire system. In contrast to periodic systems, where the current is time-dependent and only the pumped charge per pump cycle is quantized, the current itself is time independent and quantized in the quasiperiodic case [12], by the invariance of the spectrum with respect to $\varphi$.
@@
For a given pump rate $\dot{\varphi}$, the Cantor set spectrum consists of finitely many larger (adiabatic) gaps with $(\Delta E)^{2}\gg\hbar 2V\dot{\varphi}$ that are traversed adiabatically during the pumping, and infinitely many smaller (diabatic) gaps with $(\Delta E)^{2}\ll\hbar 2V\dot{\varphi}$ that are irrelevant for the pumping dynamics. Hence only the adiabatic gaps need be considered. For decreasing pump rates, more and more gaps become adiabatic and the effective bands split into new bands indexed by typically larger Chern numbers. For instance, once the 2nd-order gaps become adiabatic, the central $C=-2$ 1st-order band splits into three bands with $C=-3\,,+4\,,-3$, which together add up to $C=-2$ (see Fig. 1 c).
@@
The above division into adiabatic and diabatic gaps is particularly powerful in the regime $V/J\gg 1$, where the separation in energy gaps $\Delta E^{(n)}/\Delta E^{(n+1)}$ is largest. Moreover, in this regime, Anderson localization permits an effective model (with exponential precision in $V/J$) in which eigenstates are completely localized on a single site. In this regime, the configuration space representation offers a local, geometric view of adiabatic pumping, see Fig. 2: an increase in $\varphi$ acts as a clockwise rotation in configuration space (Fig. 2 c and d) and by continuously varying $\varphi$, every single-site localized state eventually reaches a resonance, $\theta^{(n)}$, where it resonates with another state localized $n$ sites away. The dynamics of the two resonant sites is well approximated by an effective 2-by-2 Hamiltonian
$H_{2\times 2}^{(n)}(t)=\begin{pmatrix}\delta^{(n)}(t)/2&J_{eff}^{(n)}\\ J_{eff}^{(n)}&-\delta^{(n)}(t)/2\end{pmatrix}$ (12)
@@
$-\cos{\left(\theta^{(n)}+2n\pi\beta+\varphi(t)\right)}\bigg{]}$ (13)
varies continuously, with one site going up in energy while the other goes down with $\varphi$, see inset Fig. 1a. If the gap is adiabatic ($(\Delta E)^{2}\gg\hbar 2V\dot{\varphi}$), i.e., the pump rate $\dot{\varphi}$ is slow compared to the corresponding tunneling process $J_{eff}^{(n)}\propto J^{n}/V^{n-1}$, the state adiabatically follows an eigenstate of the Rabi Hamiltonian and undergoes a Landau-Zener transition to the resonant $n$-th order neighbor – giving rise to the quantized transport. In configuration space, this Landau-Zener tunnelling translates into a discrete jump of $2\pi\beta n$.
The quantized current for a given filled band (in the case $\beta>1/2$) can be inferred as follows: In the upper first-order band, indicated by the shaded sector in Fig. 2 d, states are rotated along the blue arrow and will on average undergo a Landau-Zener transitions every time the phase winds by $\Delta\varphi=2\pi(1-\beta)$. Since in this band the transition is to the neighboring site $(n=1)$, every state is transported by on average $1/(1-\beta)$ sites per pump cycle. Combined with the fraction $(1-\beta)$ of states in the this band, the total pumped charge per pump cycle in this band is $C_{p}=(1-\beta)\frac{1}{1-\beta}=+1$, from which we extract $C=+1$. The cases where $\beta<1/2$ can be treated by sending $\beta\rightarrow 1-\beta$.
The Chern number of any given band can similarly be extracted from this local picture by computing the total current over one cycle for all states in some filled band $B^{(m,n)}$ containing only diabatically small gaps enclosed between two adiabatically large spectral gaps $\Delta E_{m},\Delta E_{n}$.
@@
Figure 3: Stability and breakdown of quantized currents under on-site disorder: Left side shows results for the lower $C=+1$ first-order band, while the right side concerns the $C=+4$ second-order band. (a,f) Thouless pumping of a particle in the clean case $W=0$, resulting in a quantized average displacement per cycle. (b,g) the particle stays confined within its energy band surrounded by open gaps throughout the Thouless pump (blue curve shows the projection of the wavefunction onto the instantaneous eigenstates in black). (c,h) In the disordered case, the averaged displacement remains quantized, despite the energy gaps being closed (d,i) by the disorder. (e,j) Average displacement per pump cycle after $2$ (e) or $1$ (j) pump periods, showing quantisation also past the closing of the clean energy gaps (red region). The observed critical disorder strengths are in excellent agreement with our proposed geometric argument (Eq. 24, Eq. 31, vertical black dotted lines), and for $C=+4$ is more than an order of magnitude larger than the gaps. Horizontal dotted lines indicate the current expected in the clean limit for an infinite pump period. Each point is obtained after averaging over $1000$ disorder realisations. Parameters: $\beta=\sqrt{2}/2$, $J=1$, $V=7.5$, $\dot{\varphi}=5\cdot 10^{-2}$ for $C=+1$, and $\dot{\varphi}=10^{-4}$ for $C=+4$.

### v3

@@ Abstract
The robustness of topological properties, such as quantized currents, generally depends on the existence of gaps surrounding the relevant energy levels or on symmetry-forbidden transitions. Here, we observe quantized currents that survive the addition of bounded local disorder beyond the closing of the relevant instantaneous energy gaps in a driven Aubry-André-Harper chain, a prototypical model of quasiperiodic systems. We explain the robustness using a local picture in configuration-space based on Landau-Zener transitions, which rests on the Anderson localisation of the eigenstates. Moreover, we propose a protocol, directly realizable in for instance cold atoms or photonic experiments, that leverages this stability to prepare topological many-body states with high Chern numbers and opens new experimental avenues for the study of both the integer and fractional quantum Hall effects.
@@
A key feature of topological phases is the robustness to local perturbations that leads to quantized physical observables characterized by a bulk topological invariant, a famous example being the integer quantum Hall effect [1, 2]. Topological phases can also be realised in dynamical systems, where time is interpreted as an additional virtual dimension. The canonical example is Thouless pumping [3, 4], where a slow time-periodic modulation of a typically 1D Hamiltonian generates quantized charge transport indexed by virtual 2D Chern numbers [5, 6]. Experimentally, Thouless pumps have been realized within various platforms including photonic waveguides [7] and ultracold atoms [8, 9, 10, 11]. They are typically considered in the context of periodic systems, where the pumped charge is only quantized when the pumping time is an exact multiple of the pump period. In contrast to periodic systems, however, Thouless pumping in quasiperiodic systems [5, 12, 13] can exhibit constant quantized pumped charge and currents at all times during the pump cycle [12].
@@
The stability of quantized Thouless pumping to disorder remains a topic of active research [4, 14, 15]. It has long been proven that the adiabatic theorem does not require a spectral gap [16], even in open quantum systems [17]. For experimentally accessible observables such as the pumped charge, however, topological quantization typically remains stable to generic perturbations only up to the closing of spectral or mobility gaps between delocalized eigenstates [18] or the breaking of explicit symmetry protections [19, 20, 21, 22]. In particular, it was shown that Thouless pumps in periodic systems break down under closing of the energy gaps by disorder [23, 24, 25].
In this letter, we report on an experimentally accessible Thouless pump for quasiperiodic systems characterized by persistent quantized pumping in the presence of large bounded local disorder, surviving the disorder-induced closure of spectral gaps. We discuss how the robustness can be readily probed within existing cold atoms experiments, and leveraged to prepare many-body states with arbitrary Chern number.
@@
To explain this counter-intuitive robustness, we develop a perturbative argument based on local Landau-Zener (avoided-crossing) transitions, and derive approximate critical disorder strengths using a configuration-space picture. The enhanced stability results from two special properties of 1D quasiperiodic Hamiltonians: First, its eigenstates can be Anderson localized [26] even in the absence of additional disorder. Second, the spectra of these quasiperiodic operators form bounded Cantor sets. Each spectrum contain a fractal hierarchy of gaps [27] whose integrated density of states (IDoS) is quantized by the gap labeling theorem [28, 29, 30], giving rise to their unusual topological properties.
@@
While our results generalize to other short-range, bounded, analytic quasiperiodic Hamiltonians, where eigenstates are Anderson localized and the gap labelling theorem holds [28, 26], we focus on the prototypical Aubry-Andre-Harper (AAH) model. This model is realized experimentally in, for instance, optical lattices [31, 32], photonic systems [5], or with cavity polaritons [33] and is known as the almost Mathieu operator in the mathematical physics literature
$\hat{H}_{AAH}^{J,V}(\varphi)=\hat{V}(\varphi)-\hat{J}.$ (1)
It consists of a nearest-neighbor hopping term
$\hat{J}=\sum_{i}J(\hat{a}^{\dagger}_{i+1}\hat{a}_{i}+\hat{a}^{\dagger}_{i}\hat{a}_{i+1})$ (2)
and an onsite cosine potential
$\hat{V}(\varphi)=\sum_{i}2V\cos(2\pi\beta i+\varphi)\hat{a}^{\dagger}_{i}\hat{a}_{i},$ (3)
@@
Figure 1: Perturbative derivation of the AAH energy spectrum: (a) AAH chain where grey line denotes the quasiperiodic modulation of onsite energies, and circled dots are lattice sites. Solid lines indicate sites connected by the tunneling $\hat{J}$. Resonances between $n$-th nearest neighbors (colored ellipses, blue for $n=1$) are mediated by an exponentially decreasing effective tunneling rate $J_{eff}^{(n)}\propto J^{n}/V^{n-1}$. (b) We map the AAH chain onto a uniformly dense unit circle referred to as configuration space. In this representation, $n$-th order neighbors are separated by an angle $\theta_{i+n}-\theta_{i}=n2\pi\beta$. Vertical direction denotes energy, and the resonances occur at fixed angles $\theta^{(n)}$, where two $n$-th order neighbors have the same height. Blue lines connect sites involved in $n=1$ resonances. (c) The resonances result in a dense set of infinitely many pairs of energy gaps labelled by the integer $n$, which for $V\gg J$ lie at energies $E^{(\pm n)}$ (we mark gaps up to $n=3$, but all gaps are open for all values of $V/J$). Sets of states surrounded by two selected gaps acquire a Chern number $C$ fixed by the labels of the surrounding gaps. Parameters: $\beta=\sqrt{2}/2$.
@@ I AAH: Energy gaps and configuration space
In periodic systems, the Brillouin zone offers a suitable base manifold on which to calculate topological invariants and band structures. To circumvent the lack of a Brillouin zone for quasiperiodic lattices, we construct a configuration-space representation of the clean AAH chain, see Fig. 1b. This representation offers a powerful language to discuss such quasiperiodic systems and enables a simple perturbative argument to derive the positions of the energy gaps, their labels, and exact expressions for their respective integrated densities of state. It is based on the observation that the only degree of freedom that differentiates one lattice site from another is the local value of the phase $\theta$ of the quasiperiodic modulation in Eq. 3. Therefore, labeling the sites with the local phase of the modulation offers a convenient description of the problem. In this configuration space representation, the position of site $i$ is mapped to the angle $\theta_{i}$ on the unit circle according to
@@
see Fig. 1b. Since $\beta$ is irrational, this map is ergodic and the set of all possible $\theta_{i}$, $\{\theta_{i}\}_{i\in\mathbb{Z}}$, densely and uniformly populates the unit circle [36], implying that every site is identified by a unique value of $\theta$.
@@
is dense on the interval $\left[-2V,2V\right]$ (see right-hand side of Fig. 1 c), and every eigenstate is localized on a single lattice site. Turning on a small tunneling amplitude ($J/V\ll 1$), resonant neighbors, i.e. nearest neighbors whose onsite energy difference is small $\delta E=2V\left|\cos{(2\pi\beta(i+1)+\varphi)}-\cos{(2\pi\beta i+\varphi)}\right|\ll J$ (marked in blue in Fig. 1 a) hybridize and form (anti-) symmetric dimers, separated by the energy $\Delta E^{(1)}=2J$. Nearly-resonant neighbors with similar onsite energies will be affected similarly.
@@
which cause the opening of pairs of gaps of size $\Delta E^{(n)}\propto 2J^{n}/V^{n-1}$ at energies
@@
Here, $[n\beta]$ refers to the integer part of $n\beta$. Eq. 9 shows that the AAH spectrum contains an infinite hierarchy of pairs of gaps – forming a dense set – labelled by the integer Chern numbers ±$\pm n$, where the two branches of the equation ensure consistency between the gap labels and the direction of quantized currents (see Appendix A for details). By the gap labeling theorem, every band, that is the set containing all states between two selected energy gaps, is indexed by a virtual Chern number [6] fixed by the difference between the two surrounding gap labels (Fig. 1 c). This Chern number is referred to as “virtual” in the sense that it is defined by treating $\varphi$ as a virtual dimension of the system. It arises from the non-commutativity of $x,\partial_{\varphi}$-translations in configuration space analogous to the non-commutativity of $x,y$-translations responsible for the integer quantum hall effect [6, 37, 38].
While the above gap construction relies on the hierarchy of effective tunnelings $J_{\textit{eff}}^{(n+1)}\ll J_{\textit{eff}}^{(n)}$ to generate a hierarchy of spectral gaps, it was shown that the spectrum indeed forms a nowhere-dense Cantor set with constant IDoS for all $V/J\neq 1$ [27, 29], extending the results of the above analysis to any value of $V/J>1$ and by self-duality to $V/J<1$.
@@ II Thouless Pumping in the AAH model
The Hamiltonian $\hat{H}^{J,V}_{AAH}(\varphi)$ is smooth and periodic in $\varphi$ and its spectrum forms a Cantor set independent of $\varphi$. Since a Cantor set is nowhere dense, the spectrum can be separated into arbitrarily many “bands”
@@
isolated by spectral gaps $\Delta E^{(m)},\Delta E^{(n)}$ at energies $E^{(m)},E^{(n)}$ independent of $\varphi\in[0,2\pi)$. This guarantees the existence of smooth contours $\gamma_{m,n}(\varphi)$ in the complex energy plane containing all of $B^{(m,n)}$ for all $\varphi$, which in turn guarantees adiabatic transport in the $\dot{\varphi}\rightarrow 0$ limit [39]. Thus, the 1D system evolving under $\hat{H}^{J,V}_{AAH}(\varphi(t))$ with $\varphi(t)=\dot{\varphi}t+\varphi_{0}$ and $(\hbar 2V\dot{\varphi})^{1/2}\ll\min(\Delta E^{(m)},\Delta E^{(n)})$ constitutes a Thouless pump with period $T=2\pi/\dot{\varphi}$ and results in a quantized current proportional to the band’s Chern number. The current
$\hat{I}(t)=\frac{i}{L}J\sum_{i=1}^{L}(\hat{a}^{\dagger}_{i+1}\hat{a}_{i}-\hat{a}^{\dagger}_{i}\hat{a}_{i+1})$ (11)
@@
is the spatial average of the current operator over the system (of length $L$), and is equal to the derivative of the total pumped charge $I\equiv\partial_{\varphi}Q$ averaged over the entire system. In our numerics, we extract the pumped charge by fitting the evolution of the center-of-mass during pumping. In contrast to periodic systems, where the current is time-dependent and only the pumped charge per pump cycle is quantized, in the quasiperiodic case the current itself is time independent and quantized [12] due to the invariance of the spectrum with respect to $\varphi$.
@@
For a given pump rate $\dot{\varphi}$, the Cantor set spectrum can be approximated to consist of finitely many larger (adiabatic) gaps with $(\Delta E)^{2}\gg\hbar 2V\dot{\varphi}$ that are traversed adiabatically during the pumping, and infinitely many smaller (diabatic) gaps with $(\Delta E)^{2}\ll\hbar 2V\dot{\varphi}$ that are irrelevant for the pumping dynamics. Hence only the adiabatic gaps need be considered. For decreasing pump rates, more and more gaps become adiabatic and the effective bands split into new bands indexed by typically larger Chern numbers. For instance, once the 2nd-order gaps become adiabatic, the central $C=-2$ 1st-order band splits into three bands with $C=-3\,,+4\,,-3$, which together add up to $C=-2$ (see Fig. 1 c).
@@
The above division into adiabatic and diabatic gaps is particularly powerful in the regime $V/J\gg 1$, where the separation in energy gaps $\Delta E^{(n)}/\Delta E^{(n+1)}$ is largest. Moreover, in this regime, Anderson localization permits an effective model (with exponential precision in $V/J$) in which eigenstates are completely localized on a single site. In this regime, the configuration space representation offers a convenient geometric view of adiabatic pumping, see Fig. 2: an increase in $\varphi$ acts as a clockwise rotation in configuration space (Fig. 2 c and d) and by continuously varying $\varphi$, every single-site localized state eventually reaches a resonance, $\theta^{(n)}$, where it resonates with another state localized $n$ sites away. The dynamics of the two resonant sites is well approximated by an effective 2-by-2 Hamiltonian
$H_{2\times 2}^{(n)}(t)=\begin{pmatrix}\delta^{(n)}(t)/2&J_{eff}^{(n)}\\ J_{eff}^{(n)}&-\delta^{(n)}(t)/2\end{pmatrix},$ (12)
@@
$-\cos{\left(\theta^{(n)}+n2\pi\beta+\varphi(t)\right)}\bigg{]}$ (13)
varies continuously, with one site going up in energy while the other goes down with $\varphi$, see the inset Fig. 1a. If the gap is adiabatic ($(\Delta E)^{2}\gg\hbar 2V\dot{\varphi}$), i.e., the pump rate $\dot{\varphi}$ is slow compared to the corresponding tunneling process $J_{eff}^{(n)}\propto J^{n}/V^{n-1}$, the state adiabatically follows an eigenstate of $H_{2\times 2}^{(n)}$ and undergoes a Landau-Zener transition to the resonant $n$-th order neighbor – giving rise to the quantized transport. In configuration space, this Landau-Zener tunnelling translates into a discrete jump of $n2\pi\beta$.
The quantized current for a given filled band (in the case $\beta>1/2$) can be inferred as follows: In the upper first-order band, indicated by the shaded sector in Fig. 2 d, states are rotated along the blue arrow and will undergo a Landau-Zener transitions every time the phase winds by $\Delta\varphi=2\pi(1-\beta)$. Since in this band the transition is to the neighboring site $(n=1)$, every state is transported by on average $1/(1-\beta)$ sites per pump cycle. Combined with the fraction $(1-\beta)$ of states in this band, the total pumped charge per pump cycle in this band is $C_{p}=(1-\beta)\frac{1}{1-\beta}=+1$, from which we extract $C=+1$. The cases where $\beta<1/2$ can be treated by sending $\beta\rightarrow 1-\beta$. The Chern number of any other filled band $B^{(m,n)}$ can be extracted analogously from this local picture by assuming that the band contains only diabatically small gaps and is enclosed by two adiabatically large spectral gaps.
Finally, this configuration-space picture directly confirms that the Chern number (gap label) of an $n$-th order gap $\Delta E^{(n)}$, i.e. the quantised current arising from filling all states up to this gap, must be equal to ±$\pm n$, since the fraction of states below a given gap is exactly equal to the phase $\Delta\varphi$ between consecutive Landau-Zener transitions, which will transport the state by $n$ sites. For determining the sign we refer the reader to Appendix A.
@@
Figure 3: Stability and breakdown of quantized currents under onsite disorder: Left side shows results for the lower $C=+1$ first-order band, while the right side depicts the $C=+4$ second-order band. (a,f) Thouless pumping of a particle in the clean case $W=0$, resulting in a quantized average displacement per cycle. (b,g) The particle stays confined within its energy band surrounded by open gaps throughout the Thouless pump. Blue curve shows the projection of the wavefunction onto the instantaneous eigenstates plotted in black. (c,h) In the disordered case, the averaged displacement remains quantized, despite the energy gaps being closed (d,i) by the disorder. (e,j) Average displacement per pump cycle after $2$ (e) or $1$ (j) pump periods, showing quantisation also past the closing of the clean energy gaps (blue regions). The critical disorder strength for $C=+4$ is more than an order of magnitude larger than the clean gaps. Horizontal dotted lines indicate the current expected in the clean limit for an infinite pump period. Each point is obtained after averaging over $1000$ disorder realisations. Parameters: $\beta=\sqrt{2}/2$, $J=1$, $V=7.5$, $\dot{\varphi}=5\cdot 10^{-2}$ for $C=+1$, and $\dot{\varphi}=10^{-4}$ for $C=+4$.